# Dyn-HTE: High-temperature expansion of the dynamic Matsubara spin correlator

### v1

@@ IV.4 Kernel trick for imaginary-time integrals
The $n+2$-fold time integral in Eq. (19) can be seen as a $n+2$ temporal dimensional Fourier transform of a time-ordered $n+2$-point correlator, but with $n$ Matsubara frequencies set to zero. Recently, in Ref. 22 closed-form analytic expressions for these Fourier transforms have been found for arbitrary $n$, generalizing the Lehmann- (or spectral-) representation beyond 2-point functions [12] (see also [52] for an earlier solution of the nested time integrals derived in the context of perturbation theory for the Hubbard model). With the help of these “kernel functions”, the Fourier transform of a time-ordered $n+2$-point correlator is written exactly in terms of many-body eigenstates, eigenenergies and matrix elements of the operators involved in the correlation function [22].
In a general case the Hamiltonian cannot be diagonalized and the “kernel trick” is typically not practically applicable. However, in the context of perturbation theory for spins, as in Eq. (19), the quantum system of interest is a set of $N$ non-interacting spins ($H_{0}=0$) with trivial product eigenstates and all eigenenergies vanishing. This allows for a tremendous simplification of the analytical expression for Eq. (19) obtained from the kernel trick. For details about the calculation, we refer to our recent work on a complementary spin-diagrammatic scheme tailored for spin systems with close-to-mean-field physics in Ref. 24. We obtain our main exact and compact result
@@
where the index ordering operator $\mathcal{P}$ sorts the operator string according to increasing superscript index in square-brackets. For example, in the case $n=2$ it acts as $\mathcal{P}V_{e_{1}}^{[4]}V_{e_{2}}^{[1]}S_{j}^{z[3]}S_{j^{\prime}}^{z[2]}=V_{e_{2}}S_{j^{\prime}}^{z}S_{j}^{z}V_{e_{1}}$.
The dimensionless kernel function $\tilde{K}_{n+2}(p_{n+1},p_{n+2};m)$ carries the frequency dependence via the Matsubara integer $m$ and depends on the positions $p_{n+1}$, $p_{n+2}$ of the external operators in the operator string. It relates to the general Kernel function $K_{n+2}$ of Ref. 22 with $n+2$ complex arguments $\Omega_{1},...,\Omega_{n+2}$ as
$\tilde{K}_{n+2}\left(a_{+},a_{-};m\right)\equiv T^{n+1}K_{n+2}(0,...,0,\underset{\mathrm{pos.}a_{+}}{\underbrace{+i\nu_{m}}},0,...,0,\underset{\mathrm{pos.}a_{-}}{\underbrace{-i\nu_{m}}},0,...,0),$ (26)
where $a_{\pm}=1,2,...,n+2$ mark the position of the frequency entry $i\nu_{\pm m}$ (of arbitrary relative order $a_{+}\gtrless a_{-})$. In Ref. 24, we showed that
@@
required by the definition (26). We remark that a somewhat similar treatment for a high-temperature expansion in the context of the fermionic Hubbard model (expansion in $t/T$ with $t$ the hopping-integral) has been put forward in Ref. 47 but to lower orders and also without the geometric flexibility of the graph-based approach.
@@ IV.5 Computational aspects for “full” graph evaluations
We finally collect a few algorithmic tricks for efficient numerical evaluation of Eq. (25) which turns out to be the performance bottleneck in the calculation of the complete graph evaluation (IV.3). This can be skipped by the reader unless interested in details of the numerical implementation or the reason for the absence of odd $l$ in the sum (17).
@@
The kernels can be precomputed for any combination of $a,\delta a$ and are only multiplied with the accumulated traces in the very end. In addition, the primed sum $\sum_{p\in S_{n}}^{\prime}$ excludes equivalent permutations of the multi-set $\{e_{1},...,e_{n}\}$ which in Eq. (25) were canceled by the factor $d[g_{n}]$.
@@
Flavor sum: So far we have not considered the structure of the edge operators which according to Sec. II are Heisenberg exchange interactions, $V_{e_{k}}=\sum_{\gamma_{k}}S_{e_{k}(1)}^{\gamma_{k}}S_{e_{k}(2)}^{\bar{\gamma}_{k}}$. Here the edge-flavor sum is over $\gamma_{k}\in\{+,-,z\}\equiv\{+1,-1,0\}$ and $e_{k}(1)$ and $e_{k}(2)$ are the two vertices connected by edge $e_{k}$. As Eq. (34) involves $n$ of these edge operators, resolving the $V_{e_{k}}$ for $k=1,...,n$ naively gives rise to $3^{n}$ flavor combinations to be summed over. As mentioned below Eq. (11), the traces $\left\langle...\right\rangle_{0}$ of strings of equal-time spin operators $\in\{S^{+}_{i},S^{-}_{i},S^{z}_{i}\}$ appearing in Eq. (34) factorize into traces of spin operators with the same vertex (site) index. The required on-site free spin equal-time correlators for a string of $n^{\prime}\leq n+2$ spin operators are pre-computed and saved as a computationally efficient integer after multiplying by a factor $1/(2S+1)^{n^{\prime}}$. The product of these factors is to be canceled by global factor $1/(2S+1)^{2n+2}$ which is to be multiplied only at the end.
It turns out that the single free spin $n$-point correlators vanish for many operator strings. For example in the case $S=1/2$ and for spin-operator strings of length $n^{\prime}=12$, only $8172$ out of the $5314413^{12}=$ 531441 possible strings have a non-zero trace. Conditions for a zero trace that do not depend on the operator order are a breaking of $U(1)$ spin rotation symmetry (unequal numbers of $S^{+}$ and $S^{-}$ in the string) or an odd number of $S^{z}$ operators which breaks the combination of time-reversal symmetry and the symmetry corresponding to a $\pi$-rotation around the spin-z axis. Such flavor combinations can be skipped already before entering the $\sum_{p\in S_{n}}^{\prime}$ sum. We also use the $\pi_{x}$-rotation symmetry in spin space that ensures equality of the traces for flavor strings connected by a global exchange of $S^{+}\leftrightarrow S^{-}$ operators.
@@ V.1 Tests of expansion coefficients
In a first step, we check the rational coefficients in the polynomials $p^{(n)}_{ii^{\prime}}(x)$ of Eq. (7) obtained from Dyn-HTE in two settings where they are known by other means, see Appendix A for details. In a first test, we consider the Heisenberg four spin cluster ($N=4$) with all-to-all interactions for which the Matsubara correlator $G_{ii^{\prime}}(i\nu_{m})$ can be found analytically for $S=1/2$, $S=1$ and for the local and non-local case. The rational coefficients of a series expansion of these exact results in $x=J/T$ agrees to our graph-embedding based Dyn-HTE result. This test thus involves evaluations for all graphs with up to four vertices (they are embeddable in the $N=4$ all-to-all cluster). In a second test we focus on extended lattices in two and three dimensions (like kagome, triangular, pyrochlore, …) and compute their Dyn-HTE. At each order in $x$, we take the frequency- and site-sums to reproduce the conventional HTE expansion of the (purely static) uniform susceptibility $\chi$ found in the conventional HTE literature.
@@ V.2 Benchmark: Heisenberg S=1/2 AFM chain
The summation of the bare Dyn-HTE series up to order $n=12$ (gray line) starts to deviate significantly from the exact results at rather high temperature $x=J/T\simeq 2$. This is in close analogy to conventional HTE where it is well known to signal the small radius of convergence of the $x$-series in the complex plane [2]. A standard tool to extract meaningful information from the bare HTE series at smaller temperatures are Padé approximants [2], which are rational functions ${[K,L]}(x)=\frac{P_{K}(x)}{Q_{L}(x)}$ with polynomials $P_{K}$ and $Q_{L}$ of degree $K$ and $L$, respectively. For $K+L\leq n_{max}$, their coefficients can be determined so that the series expansion of the Padé approximant agrees in the first $n_{max}$ orders with the bare (Dyn-)HTE series. For the case $[K=6,L=6]$ [blue lines in Fig. 3], the agreement with QMC data already extends to $x\simeq 4$, both for the real- and momentum-space correlator. Note that for the latter, we perform the Fourier transform before computing the Padé approximant.
For many combinations of Matsubara integer $m$ and distance $i-i^{\prime}$ (or momentum $k$) the [6,6] Padé approximant agrees with the error controlled QMC result to even much larger $x\simeq 8$, most severe deviations occur for the $m=0$ case with $i-i^{\prime}=1$ or $k=\pi$. In real application without benchmark data, the quality of a prediction from Padé approximants is usually gauged by comparison of different $[K,L]$ with $K+L\leq n_{max}$ [2]. Due to the asymptotically constant (and finite) values of $TG_{ii^{\prime}}(i\nu_{m})$ at $x\rightarrow\infty$, we limit ourselves to the comparison of the symmetric Padé approximants [6,6] and [5,5] (dashed lines). Indeed, the poor quality of the [6,6] approximants in the two cases mentioned above is reflected in a large disagreement between the two Padé approximants [5,5] and [6,6].
@@ V.4 Renormalized mean-field form of static susceptibility
In Ref. 58 we could show analytically that the HTE of the beyond-rMF terms $\epsilon_{n}$ for $n=2,3,...$ start at order $O(x^{4})$ with a particularly small prefactor. This lead us to conjecture the excellent validity of the rMF beyond the high-temperature regime which was corroborated by matching the rMF form (37) with published momentum-dependent susceptibilities obtained from the BDMC method [31] and pseudo-fermion functional renormalization group [59].
@@
Figure 5: Amplitudes of parameter-ratios of the rMF form (37): $x/f_{x}$, $g_{x}/f_{x}$ and the first few corrections $\epsilon_{2,3,4}/f_{x}$ [c.f. Eq. (36)] for the nearest-neighbor $S=1/2$ Heisenberg AFM model on the chain, triangular, kagome and pyrochlore lattice from Dyn-HTE. For resummation, we used $[6,6]$ and $[5,5]$ u-Padé approximants. The data is multiplied by the maximum (eigenvalue) of $-\gamma_{1}(\mathbf{k})$ given in the panels alongside the parameters $f$. For the chain geometry QMC data (dots) [54] matches well with the Dyn-HTE data (lines). For triangular, kagome and pyrochlore lattice $\epsilon_{2,3,4}/f_{x}$ are of order $10^{-3}$ rendering the rMF an excellent approximation. Whenever the [6,6] or [5,5] Padé approximant is anomalous (with obvious poles), we use a stable lower-order Padé approximant.
A quantitative measure of the rMF form’s accuracy is provided by the ratio $|\frac{\sum_{n}\epsilon_{n}\gamma_{n}(\mathbf{k})}{f_{x}+g_{x}\gamma_{1}(\mathbf{k})}|$ for which values much smaller than unity indicate the applicability of the approximation. In Fig. 5 we report several ratios of $f_{x}$, $g_{x}$ and the numerically dominant $\epsilon_{2},\epsilon_{3},\epsilon_{4}$ for the Heisenberg $S=1/2$ AFM on various lattices. They are obtained from the $u$-Padé approximations for the $x$-series expansions of the respective ratio found from Dyn-HTE. For the chain (top left panel), the ratios $1/f_{x}$, $g_{x}/f_{x}$ and $\epsilon_{n}/f_{x}$ saturate for low $T$ (large $x$), hence resummation of the ratios with the Padé approximant of the $u$-series works very well as witnessed by comparison to QMC (dots). As $\epsilon_{2}$ grows sizably for low $T$, the validity of the rMF approximation is compromised. This was stated before in Ref. 60.
@@ VI Conclusion and outlook
We have presented a dynamical extension of the high-temperature series expansion (HTE) to the Matsubara spin correlator for spin Hamiltonians (Dyn-HTE). Currently, our numerical implementation [21] features an expansion to order $n_{max}=12$ and is applicable for $S\leq 1$ Heisenberg Hamiltonians with a single coupling constant $J$ on arbitrary (in particular frustrated and high-dimensional) lattices. The real-frequency dynamic structure factor can also be obtained via postprocessing of the Matsubara data as described in our companion letter [34].
For further methodological development of Dyn-HTE it would be worthwhile to investigate the potential of the linked-cluster method [2]. Here, only graphs (clusters) with simple edges are required on which the observable needs to be expanded in $x=J/T$ or is even determined by exact diagonalization. While this could be a viable way to higher expansion orders for small $S$, it would also be interesting to extend Dyn-HTE in the form presented in this work to arbitrary $S$. This could be achieved by building on closed-form expressions for the equal-time free spin correlators [22].
Moreover, possible extensions of Dyn-HTE analogous to achievements in conventional HTE include the application to Heisenberg models with more than one coupling constant, e.g. $J_{1}$-$J_{2}$ models [5], single-ion anisotropies or magnetic fields [6]. In the latter case, the kernel trick for the evaluation of the imaginary-time integrals needs to be adapted for the presence of (a few) non-zero many-body eigenenergies of $H_{0}$.
Beyond the Heisenberg case, Dyn-HTE can be extended to models with broken or reduced spin rotation symmetry like the XXZ case. Likewise, also the fermionic (or bosonic) $t$-$J$-model which is prominently realized in cold-atom quantum simulation [61] should be considered.
Finally, it would be interesting to aim the Dyn-HTE to more complex correlation functions (e.g. for higher-order response) or improve the series convergence beyond the standard Padé approximants by implementing advanced ideas like, e.g., the homotopic action [62].

### v2

@@ IV.4 Kernel trick for imaginary-time integrals
The central quantity left to be computed is $c_{g^{(n)}}^{[full]}(i\nu_{m})$ in Eq. (19), the full part of the graph evaluation. Its $n+2$-fold time integral can be seen as a $n+2$ dimensional temporal Fourier transform of an (imaginary-)time-ordered $n+2$-point correlator, but with $n$ Matsubara frequencies set to zero. Recently, in Ref. 22 closed-form analytic expressions for these Fourier transforms have been found for arbitrary $n$, generalizing the Lehmann representation [12] beyond 2-point functions [50] (see also [51] for an earlier solution of the nested time integrals derived in the context of perturbation theory for the Hubbard model). With the help of these “kernel functions”, the Fourier transform of a time-ordered $n+2$-point correlator is written exactly in terms of many-body eigenstates, eigenenergies and matrix elements of the operators involved in the correlation function [22].
In a general case the Hamiltonian cannot be diagonalized and the “kernel trick” is typically not practically applicable. However, in the context of spins, as in Eq. (19), the quantum system of interest is a set of $N$ non-interacting spins ($H_{0}=0$) with trivial product eigenstates and all eigenenergies vanishing. This allows for a tremendous simplification of the analytical expression for Eq. (19) obtained from the kernel trick. For details about the calculation, we refer to our recent work on a complementary spin-diagrammatic scheme tailored for spin systems with close-to-mean-field physics in Ref. 52.
We obtain our main exact and compact result
@@
where the index ordering operator $\mathcal{P}$ sorts the operator string according to increasing superscript index (in square-brackets). For example, in the case $n=2$ and for a particular permutation $p$ it acts as $\mathcal{P}V_{e_{1}}^{[4]}V_{e_{2}}^{[1]}S_{j}^{z[3]}S_{j^{\prime}}^{z[2]}=V_{e_{2}}S_{j^{\prime}}^{z}S_{j}^{z}V_{e_{1}}$.
The dimensionless kernel function $\tilde{K}_{n+2}(p_{n+1},p_{n+2};m)$ carries the dependence on external frequency via the Matsubara integer $m$ and depends on the positions $p_{n+1}$, $p_{n+2}$ of the external operators in the operator string. It relates to the general Kernel function $K_{n+2}$ of Ref. 22 with $n+2$ complex arguments $\Omega_{1},...,\Omega_{n+2}$ as
$\tilde{K}_{n+2}\left(a_{+},a_{-};m\right)\equiv T^{n+2}K_{n+2}(0,...,0,\underset{\mathrm{pos.}a_{+}}{\underbrace{+i\nu_{m}}},0,...,0,\underset{\mathrm{pos.}a_{-}}{\underbrace{-i\nu_{m}}},0,...,0),$ (25)
where $a_{\pm}=1,2,...,n+2$ mark the position of the frequency entry $i\nu_{\pm m}$ (of arbitrary relative order $a_{+}\gtrless a_{-})$. In Ref. 52, we showed that
@@
required by the definition (25). We remark that a somewhat similar treatment for a high-temperature expansion in the context of the fermionic Hubbard model (expansion in $t/T$ with $t$ the hopping-integral) has been put forward in Ref. 44 but to lower orders and also without the geometric flexibility of the graph-based approach.
In passing, we note a variety of recently suggested approaches that have been developed to solve Matsubara sums (or, equivalently, imaginary-time integrals) that appear in bare or renormalized perturbation theory analytically. These include the discrete-Lehmann representation [53], the intermediate representation [54] and algorithmic Matsubara integration [55, 56]. However, none of these approaches are applicable to perturbation theory diagrams with general bosonic $n$-point correlators as building blocks, like in the present expansion (which can be understood to be of strong-coupling type in Hubbard-model parlance).
@@ IV.5 Computational aspects for “full” graph evaluations
We finally collect a few algorithmic tricks for efficient numerical evaluation of Eq. (24) which turns out to be the performance bottleneck in the calculation of the complete graph evaluation (IV.3). This can be skipped by the reader unless interested in details of the numerical implementation or the reason for the absence of odd $l$ in the sum (17). With these tricks in place, the evaluations for all graphs up to and including order $n_{max}=12$ required on the order of $500.000$ core hours for spin length $S=1/2$.
@@
The kernels can be precomputed for any combination of $a,\delta a$ and are only multiplied with the accumulated traces ($\sim\left\langle...\right\rangle_{0}$) in the very end. In addition, the primed sum $\sum_{p\in S_{n}}^{\prime}$ excludes equivalent permutations of the multi-set $\{e_{1},...,e_{n}\}$ which in Eq. (24) were canceled by the factor $d[g_{n}]$.
@@
Flavor sum: So far we have not considered the structure of the edge operators which according to Sec. II are Heisenberg exchange interactions, $V_{e_{k}}=\sum_{\gamma_{k}}S_{e_{k}(1)}^{\gamma_{k}}S_{e_{k}(2)}^{\bar{\gamma}_{k}}$. Here the edge-flavor sum is over $\gamma_{k}\in\{+,-,z\}\equiv\{+1,-1,0\}$ and $e_{k}(1)$ and $e_{k}(2)$ are the two vertices connected by edge $e_{k}$. As Eq. (33) involves $n$ of these edge operators, resolving the $V_{e_{k}}$ for k=1,,nk=1,...,n naively gives rise to $3^{n}$ flavor combinations to be summed over. As mentioned below Eq. (11), the traces $\left\langle...\right\rangle_{0}$ of strings of equal-time spin operators $\in\{S^{+}_{i},S^{-}_{i},S^{z}_{i}\}$ appearing in Eq. (33) factorize into traces of spin operators with the same vertex (site) index. The required on-site free spin equal-time correlators for a string of $n^{\prime}\leq n+2$ spin operators are pre-computed and saved as a computationally efficient integer after multiplying by a factor $(2S+1)^{n^{\prime}}$. The product of these factors is to be canceled by global factor $1/(2S+1)^{2n+2}$ which is to be multiplied only at the end.
It turns out that the single free spin $n$-point correlators vanish for many operator strings. For example in the case $S=1/2$ and for spin-operator strings of length $n^{\prime}=12$, only $8172$ out of the $3^{12}=531441$ possible strings have a non-zero trace. Conditions for a zero trace that do not depend on the operator order are a breaking of $U(1)$ spin rotation symmetry (unequal numbers of $S^{+}$ and $S^{-}$ in the string) or a string that is odd under a $\pi$-rotation around the spin-x axis. The latter is the case for a string consisting of an odd number of $S^{z}$ operators, $S^{z}S^{z}...S^{z}$. Such flavor combinations can be skipped already before entering the $\sum_{p\in S_{n}}^{\prime}$ sum. We also use the $\pi_{x}$-rotation symmetry in spin space that ensures equality of the traces for flavor strings connected by a global exchange of $S^{+}\leftrightarrow S^{-}$ operators.
@@ V.1 Tests of expansion coefficients
In a first step, we check the rational coefficients in the polynomials $p^{(n)}_{ii^{\prime}}(x)$ of Eq. (7) obtained from Dyn-HTE in two settings where they are known by other means, see Appendix A for details. In a first test, we consider the Heisenberg four spin cluster ($N=4$) with all-to-all interactions for which the Matsubara correlator $G_{ii^{\prime}}(i\nu_{m})$ can be found analytically for $S=1/2$, $S=1$ and both for the local and non-local case. The rational coefficients of a series expansion of these exact results in $x=J/T$ agrees to our graph-embedding based Dyn-HTE result. This test thus involves evaluations for all graphs with up to four vertices (as they are embeddable in the $N=4$ all-to-all cluster). In a second test we focus on extended lattices in two and three dimensions (like kagome, triangular, pyrochlore, …) and compute their Dyn-HTE. At each order in $x$, we take the frequency- and site-sums to reproduce the conventional HTE expansion of the (purely static) uniform susceptibility $\chi$ found in the conventional HTE literature.
@@ V.2 Benchmark: Heisenberg S=1/2 AFM chain
The summation of the bare Dyn-HTE series up to order $n=12$ (gray line) starts to deviate significantly from the exact results at rather high temperature $x=J/T\simeq 2$. This is in close analogy to conventional HTE where it is well known to signal the small radius of convergence of the $x$-series in the complex plane [2]. A standard tool to extract meaningful information from the bare HTE series at smaller temperatures are Padé approximants [2], which are rational functions ${[K,L]}(x)=\frac{P_{K}(x)}{Q_{L}(x)}$ with polynomials $P_{K}$ and $Q_{L}$ of degree $K$ and $L$, respectively. For $K+L=n_{max}$, their coefficients can be determined so that the series expansion of the Padé approximant agrees in the first $n_{max}$ orders with the bare (Dyn-)HTE series. For the case $K=6,L=6]\equiv[6,6]$, denoted by blue lines in Fig. 3, the agreement with QMC data already extends to $x\simeq 4$, both for the real- and momentum-space correlator. Note that for the latter, we perform the Fourier transform before computing the Padé approximant.
For many combinations of Matsubara integer $m$ and distance $i-i^{\prime}$ (or momentum $k$) the [6,6] Padé approximant agrees with the error controlled QMC result to even much larger $x\simeq 8$, most severe deviations occur for the $m=0$ case with $i-i^{\prime}=1$ or $k=\pi$. In real applications (without available benchmark data), the quality of a prediction from Padé approximants is usually gauged by comparison of different $[K,L]$ with $K+L\leq n_{max}$ [2]. Due to the asymptotically constant (and finite) values of $TG_{ii^{\prime}}(i\nu_{m})$ at $x\rightarrow\infty$, we limit ourselves to the comparison of the symmetric Padé approximants [6,6] and [5,5] (dashed lines). Indeed, the poor quality of the [6,6] approximants in the two cases mentioned above is reflected in a large disagreement between the two Padé approximants [5,5] and [6,6].
@@ V.4 Renormalized mean-field form of static susceptibility
In Ref. 62 we could show analytically that the HTE of the beyond-rMF terms $\epsilon_{n}$ for $n=2,3,$... start at order $O(x^{4})$ with a particularly small prefactor. This lead us to conjecture the excellent validity of the rMF beyond the high-temperature regime which was corroborated by matching the rMF form (36) with published momentum-dependent susceptibilities obtained from the BDMC method [30] and pseudo-fermion functional renormalization group [63].
@@
Figure 5: Amplitudes of parameter-ratios of the rMF form (36): $x/f_{x}$, $g_{x}/f_{x}$ and the first few corrections $\epsilon_{2,3,4}/f_{x}$ [c.f. Eq. (35)] for the nearest-neighbor $S=1/2$ Heisenberg AFM model on the chain, triangular, kagome and pyrochlore lattice from Dyn-HTE. For resummation, we used $[6,6]$ and $[5,5]$ u-Padé approximants. The data is multiplied by the maximum (eigenvalue) of $\gamma_{1}(\mathbf{k})$ given in the panels alongside the u-Padé parameter $f$. For the chain geometry QMC data (dots) [58] matches well with the Dyn-HTE data (lines). For triangular, kagome and pyrochlore lattice $\epsilon_{2,3,4}/f_{x}$ are of order $10^{-3}$ rendering the rMF an excellent approximation. Whenever the [6,6] or [5,5] Padé approximant is anomalous (with obvious poles), we use a stable lower-order Padé approximant.
A quantitative measure of the rMF form’s accuracy is provided by the ratio $|\frac{\sum_{n}\epsilon_{n}\gamma_{n}(\mathbf{k})}{f_{x}+g_{x}\gamma_{1}(\mathbf{k})}|$ for which values much smaller than unity indicate the applicability of the approximation. In Fig. 5 we report several ratios of $f_{x}$, $g_{x}$ and the numerically dominant $\epsilon_{2},\epsilon_{3},\epsilon_{4}$ for the Heisenberg $S=1/2$ AFM on various lattices. They are obtained from the $u$-Padé approximations for the $x$-series expansions of the respective ratio found from Dyn-HTE. For the chain (top left panel), the ratios $1/f_{x}$, $g_{x}/f_{x}$ and $\epsilon_{n}/f_{x}$ saturate for low $T$ (large $x$), hence resummation of the ratios with the Padé approximant of the $u$-series works very well as witnessed by comparison to QMC (dots). As $\epsilon_{2}$ grows sizably for low $T$, the validity of the rMF approximation is compromised. This was stated before in Ref. 64.
@@ VI Conclusion and outlook
We have presented a dynamical extension of the high-temperature series expansion (HTE) to the Matsubara spin correlator for spin Hamiltonians (Dyn-HTE). Currently, our numerical implementation [21] features an expansion to order $n_{max}=12$ and is available for $S\leq 1$ Heisenberg Hamiltonians with a single coupling constant $J$ on arbitrary (in particular frustrated and high-dimensional) lattices. The real-frequency dynamic structure factor can be obtained via postprocessing of the Matsubara data as described in our companion letter [39].
For further methodological development of Dyn-HTE it would be worthwhile to reduce the number of graphs by the free-graph expansion technique (where different vertices of a graph can be assigned to the same or different lattice sites) or by the linked-cluster method [2]. In the latter, only graphs (clusters) with simple edges are required on which the observable needs to be expanded in $x=J/T$ or is even determined by exact diagonalization. While this could be a viable way to higher expansion orders for small $S$, it would also be interesting to extend Dyn-HTE in the form presented in this work to arbitrary $S$. This could be achieved by building on closed-form expressions for the equal-time free spin correlators [22].
Moreover, possible extensions of Dyn-HTE analogous to achievements in conventional HTE include the application to Heisenberg models with more than one coupling constant, e.g. $J_{1}$-$J_{2}$ models [5], single-ion anisotropies or magnetic fields [6]. In the latter case, the kernel trick for the evaluation of the imaginary-time integrals needs to be adapted for the presence of (a few) non-zero many-body eigenenergies of $H_{0}$. It would also be interesting to aim Dyn-HTE at more complex correlation functions like three-point correlators encoding higher-order response, see Ref. [28] for pioneering work. Likewise, it would be useful to improve the series convergence beyond the standard Padé approximants by implementing advanced ideas like, e.g., the homotopic action [65].
Beyond the Heisenberg case, Dyn-HTE can be extended to models with broken or reduced spin rotation symmetry like the XXZ case. Likewise, also the fermionic (or bosonic) $t$-$J$-model which is prominently realized in cold-atom quantum simulation [66] should be considered. Here, existing treatments [67, 44] could be extended to higher expansion orders.